What is the third term of a recursive sequence that follows the pattern x - 7 and has a first term x = 100?
Found 2 solutions by rapaljer, uma:
Answer by rapaljer(4671)

If the first term is 100, the second term would be 93, and the third term would be 7 less than this which would be 86. Right???

R^2

Answer by uma(370)

First term is x = 100
So second term = x - 7
= 100 - 7
= 93
Thus the third term = 93 - 7
= 86

Good Luck!!!